# Driving in the Occupancy World: Vision-Centric 4D Occupancy Forecasting and Planning via World Models for Autonomous Driving

### v2

@@ 4.4 Ablation Study
The default configuration for the ablation experiments involves using one historical and the current images as input to predict the inflated occupancy over two future timestamps.
@@ 5 Conclusion
Acknowledgements:
We thank Jiangning Zhang and Jiang He for helpful discussions and valuable support on the paper. We thank all authors for their contributions.
@@ C.1 Dataset
We adopt the nuScenes (Caesar et al. 2020) and nuScenes-Occupancy (Wang et al. 2023c) datasets for occupancy forecasting and planning experiments, utilizing 700 scenes for training and 150 for validation. These datasets provide 16 classes of semantic occupancy annotations, from which bicycle, bus, car, construction, motorcycle, trailer, truck, and pedestrian are selected as movable objects for flow forecasting and collision avoidance during planning. Following Cam4DOcc (Ma et al. 2024b), we use the corresponding images from two historical frames and one current frame as input to predict future states over four timestamps. After reorganizing the data, the number of sequences for training and testing is 23,930 and 5,119, respectively.

### v3

@@ 4.4 Ablation Study
The default configuration for the ablation experiments involves using one historical and the current images as input to predict the inflated GMO over two future timestamps.
@@ 5 Conclusion
Acknowledgments
We thank Jiangning Zhang and Jiang He for helpful discussions and valuable support on the paper. We thank all authors for their contributions. This work was supported by a Grant from The National Natural Science Foundation of China (No. 62103363).
@@ C.1 Dataset
We adopt the nuScenes (Caesar et al. 2020), nuScenes-Occupancy (Wang et al. 2023c), and Lyft-Level5 datasets for occupancy forecasting and planning. Following Cam4DOcc (Ma et al. 2024b), we use the corresponding images from two historical frames and one current frame as input to predict future states over four timestamps. Since nuScenes is annotated at 2Hz while Lyft-Level5 is annotated at 5Hz, we report the results with 2s and 0.8s time intervals, respectively. After reorganizing the data, the number of sequences for training and testing is 23,930 and 5,119 in scenes and 15,720 and 5,880 in Lyft-Level5.